# SlowFast-SCI: Slow-Fast Deep Unfolding Learning for Spectral Compressive Imaging

### v1

@@ 3.2 Fast Learning
Our slow learning phase (Sec. 3.1) endows the deep unfolding network with robust physics‐informed priors and learnable parameters. Yet when faced with a new spectral domain, naively fine‐tuning every unfolding block demands as much data and compute as the original training. To overcome this, we observed that each denoiser—when linearized—collapses into an optimal Wiener filter qin2023ground ; zhao2024adapnet , which led us to insert lightweight adaptation blocks into each stage and self‐supervisedly fine‐tune only these per test sample (Sec. 3.2.1). Then, we distill the pre-trained network into a nimble ‘fast unfolding’ variant and finally introduce a customized self-supervised imaging guidance loss on the adapters (Sec. 3.2.2), delivering rapid inference and seamless cross-domain adaptation.
3.2.1 Fast Learning via Wiener-Filter–Inspired Adapter Modules
Let $\mathcal{D}(\cdot,\omega)$ denote a denoising network with parameters $\omega$. Then the Eq. (5) is transformed into the following form:
$\mathrm{\mathbf{x}}_{k+1}=\mathcal{D}(\mathrm{\mathbf{r}}_{k+1},\omega_{k+1}).$ (6)
Training $\mathcal{D}(\cdot,\omega)$ to minimize prediction error means minimizing the mean squared error (MSE) by updating the parameters $\omega$:
$\min_{\theta}\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x}),\mathbf{\epsilon}\sim p(\mathbf{\epsilon})}\{\|\mathcal{D}(\mathrm{\mathbf{r}},\omega)-\mathbf{x}\|_{F}^{2}\}\approx\min_{\theta}\sum_{i}\|\mathcal{D}(\mathrm{\mathbf{r}}_{(i)},\omega)-\mathbf{x}_{(i)}\|_{F}^{2}\}.$ (7)
How to understand the statistical properties of general CNNs as a form of Bayesian inference is an open question. For the sake of discussion, we consider a CNN consisting of only linear convolution layers, which does not contain nonlinear activation functions and bias terms, and its expression is:
$\mathcal{D}(\mathrm{\mathbf{r}},\omega)=\mathcal{H}\otimes\mathrm{\mathbf{r}},$ (8)
where $\otimes$ represents the convolution operator and $\mathcal{H}$ is the filter of $N_{\lambda}$ channels. Let $\mathcal{H}_{w}$ denote the optimal filter, Wiener filter, we have, for each channel $\lambda_{i}$,
$\widehat{\mathcal{H}}_{w}[k_{1},k_{2},\lambda_{i}]=\frac{\mathbb{E}_{\mathbf{x}}\{|\widehat{\mathbf{x}}[k_{1},k_{2},\lambda_{i}]|^{2}\}}{\mathbb{E}_{\mathbf{x}}\{|\widehat{\mathbf{x}}[k_{1},k_{2},\lambda_{i}]|^{2}\}+\mathbb{E}_{\mathbf{\epsilon}}\{|\widehat{\mathbf{\epsilon}}[k_{1},k_{2},\lambda_{i}]|^{2}\}},$ (9)
where $k_{1},k_{2}]$ denotes the frequency of the image in the frequency domain after discrete Fourier transformation (DFT). The derivation details of Eq. (9) are given in the supplementary material. Similarly, for a specific test sample $\tilde{\mathrm{\mathbf{r}}}=\tilde{\mathrm{\mathbf{x}}}+\tilde{\mathrm{\mathbf{\epsilon}}}$, its individual optimal wiener filter for each channel $\lambda_{i}$ is given by
@@ 3.2.1 Fast Learning via Wiener-Filter–Inspired Adapter Modules
Pre-training a linear CNN on source data effectively learns the Wiener estimator for that distribution—derived from the image and noise second‐order statistics in the frequency domain (Eq. (9)). Since the Wiener solution for a specific target scene (Eq. (10)) only differs modestly from the source estimator, a compact, supplementary convolutional filter can bridge this gap, aligning the pre-trained source filter with the target‐scene Wiener filter.
3.2.2 Self-Adaptation-Driven Fast-Learning Unfolding
Imaging-Guided Fast Unfolding Distillation. The process of model distillation is shown in Fig. 2. During model distillation, the pre-trained model, with its parameters frozen, serves as the teacher model, while the two-stage model with the same architecture as the pre-trained model is used as the student model. The dataset used for distillation is a dataset that the pre-trained model has not seen during supervised slow learning. Firstly, mask $\Phi$ and measurements $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}\in\mathbb{R}^{H\times(W+d\times(N_{\lambda}-1)}$ are entered into the teacher model to obtain the results $\{\mathrm{\mathbf{X}}_{th,i}\}_{i=1}^{N_{d}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$, which will be used as a label to supervise the training of the student model. Then mask $\Phi$ and measurements $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}$ are input into the student model to obtain the results $\{\mathrm{\mathbf{X}}_{st,i}\}_{i=1}^{N_{d}}$. We update the student model by minimizing the following distillation loss:
$\mathscr{L}_{dis}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\|\mathrm{\mathbf{X}}_{st,i}-\mathrm{\mathbf{X}}_{th,i}\|_{F}^{2}.$ (11)
The inference speed of the two-stage student model obtained by model distillation is much faster than the original pre-trained model, and its performance on test data is comparable to that of the original pre-trained model.
Adapter-Enhanced Unfolding Network Architecture. Motivated by the discussion in Sec. 3.2.1, we propose a two-stage unfolding network where each stage is attached by a single lightweight module, namely FAST-TTAM, which takes both spatial and spectral information into account. The distilled unfolding network, is a typical unfolding network for SCI reconstruction. For each stage of the distilled unfolding network, we concatenate the adjustable lightweight operator to the end of the frozen stage. The FAST-TTAM is a lightweight operator with one update step, one parameter $\mu$ estimation network, and two convolution blocks, the spatial convolution block and the spectral convolution block. The structure of parameter $\mu$ estimation network is shown in the upper right of Fig. 2, which is used to estimate the parameter $\mu$. The spatial convolution block consists of two convolution layers with $3\times 3$ filter and one $\mathrm{ReLU}$ function, while the spectral convolution block only includes one convolution layer with $1\times 1$ filter.
Self-Supervised Training and Test-Time Adaptation. Consider the distilled unfolding network with adaptive layers $\mathcal{F}_{DA}(\cdot;\omega^{*},\theta)$:
$\mathcal{F}_{DA}(\cdot;\omega^{*},\theta):\mathrm{\mathbf{y}},\mathbf{\Phi}\rightarrow{\mathrm{\mathbf{x}}},$ (12)
where $\omega^{*}$ denotes the parameters of the distilled model frozen in self-supervised training and test-time adaptation and $\theta$ denotes the parameters of the adapt layers to be updated. Given the same dataset $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}$ as in model distillation, we update $\theta$ by minimizing the following self-supervised loss:
[…]
$\mathbb{R}^{N}=\mathrm{im}(\mathbf{\Phi}^{H})\oplus\mathrm{ker}(\mathbf{\Phi}),$ (14)
where $\mathrm{im}(\mathbf{\Phi}^{H})=\{\mathbf{\Phi}^{H}\mathrm{\mathbf{y}},\mathrm{\mathbf{y}}\in\mathbb{R}^{M}\}$ and $\mathrm{ker}(\mathbf{\Phi})=\{\mathrm{\mathbf{x}}\in\mathbb{R}^{N},\mathbf{\Phi}\mathrm{\mathbf{x}}=\mathbf{0}\}$. To solve the severe linear inverse problem of SCI during test-time adaptation, the key is how to adapt the model for prediction with small prediction errors in both $\mathrm{im}(\mathbf{\Phi}^{H})$ and $\mathrm{ker}(\mathbf{\Phi})$ efficiently at the test time. To address this, we propose a test-time self-supervised loss function $\mathscr{L}_{tta}$:
$\mathscr{L}_{tta}:=\mathscr{L}_{im}+\lambda\mathscr{L}_{ker},$ (15)
with a pre-defined hyperparameter $\lambda$. $\mathscr{L}_{im}$ and $\mathscr{L}_{ker}$ concern the prediction errors in $\mathrm{im}(\mathbf{\Phi}^{H})$ and $\mathrm{ker}(\mathbf{\Phi})$ respectively. For the prediction in $\mathrm{im}(\mathbf{\Phi}^{H})$, we use the measurement $\mathrm{\mathbf{y}}$ to train our network.
$\mathbf{\Phi}\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta)\rightarrow\mathrm{\mathbf{y}}\approx\mathbf{\Phi}\mathrm{\mathbf{x}}+\mathbf{n}.$ (16)
For the prediction in $\mathrm{ker}(\mathbf{\Phi})$, we train our model using the output $\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta)$ as label and $\mathbf{\Phi}T(\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta))$ as input, where $T(\cdot)$ represents the data augmentation (shifts, rotations, etc.).
$\mathcal{F}_{DA}(\mathbf{\Phi}T(\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta);\omega^{*},\theta))\rightarrow T(\mathcal{F}_{DA}(\mathrm{\mathbf{y}};\omega^{*},\theta);\omega^{*},\theta)).$ (17)
@@ 3.2.2 Self-Adaptation-Driven Fast-Learning Unfolding
Table 1: Comparisons (PSNR (upper entry ineach cell), SSIM (lower entry in each cell), Params, and FLOPs) between SlowFast-SCI (SlowFast-SCI1 and SlowFast-SCI2 represent the pre-trained model we used were DAUHST and DPU, respectively.) and SOTA methods on 10 simulation scenes of Harvard dataset.
@@
MST mst CVPR 2022 1.92 25.60
32.67
0.857
30.28
0.699
31.20
0.798
26.21
0.515
38.63
0.938
30.40
0.837
32.80
0.751
29.34
0.791
32.40
0.732
27.85
0.670
31.18
0.759
CST cst ECCV 2022 2.86 36.77
33.17
0.882
32.29
0.789
32.90
0.845
27.74
0.595
39.44
0.960
29.29
0.871
33.60
0.812
31.53
0.867
34.00
0.811
28.60
0.733
32.26
0.817
BIRNAT bisci TPMAI 2023 4.19 1973.66
35.58
0.914
35.23
0.884
34.93
0.864
30.23
0.721
42.38
0.976
35.97
0.917
36.28
0.896
33.62
0.907
36.87
0.888
30.48
0.807
35.16
0.877
RDLUF Dong_2023_CVPR_rdluf CVPR 2023 1.81 115.34
@@
Figure 4: Simulated HSI reconstruction comparisons (middle) on Harvard. The left shows the scene measurement and the right shows the spectral curves corresponding to the selected region. Please zoom in for better view.
4 Experiments
4.1 Experiment Setup
Datasets. The pre-trained model used in slow learning was trained on the CAVE park2007multispectralcave dataset. We adopted two datasets Harvard chakrabarti2011statistics_harvard and ICVL arad2016sparse_icvl with randomly clipped patches of size $256\times 256\times 28$ for simulation experiments. Harvard/ICVL dataset contains 67/180 images for training and 10/20 for testing. For real datasets, we used 5 scenes with spatial size $660\times 660$ and 28 bands captured in TSA-Net meng2020endtsa_cassi , then cropped data blocks of size $256\times 256\times 28$ for testing.
Comparison Methods. We compare our SlowFast-SCI on synthetic data and real scenes with state-of-the-art (SOTA) reconstruction methods including TSA-Net meng2020endtsa_cassi , MSTmst , CSTcst , BIRNATbisci , RDLUFDong_2023_CVPR_rdluf , SSR-L Zhang_2024_CVPR_ssr , SAH-SCI liu2024sahsci , DAUHST dauhst , and DPU Zhang_2024_CVPR_dpu .
Model Distillation and FAST-TTAM Training. In model distillation, we distilled two typical deep unfolding networks with 9 stages, DAUHST darestani2022test_tta and DPU Zhang_2024_CVPR_dpu . The distillation epoch and batch size were set to 200 and 2, respectively. We used the Adam 2014Adam optimizer for both networks. We set the initial learning rate as $4\times{10}^{-4}$ and adopted the Cosine Annealing learning rate scheme loshchilov2016sgdr to implement training. We trained the FAST-TTAM for 100 epochs. The batch size was set to 2 and the learning rate was fixed at $1\times{10}^{-4}$. The $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$ were set to 0.7, 0.4 and 0.001 in Eq. (13).
Test-Time Adaptation. The test-time adaptation was conducted over 50 iterations for Harvard and 10 iterations for ICVL, with a constant learning rate of $1\times{10}^{-3}$ and $\lambda=0.7$ throughout all iterations.
4.2 Quantitative Results
Comparison on Harvard and ICVL.
Tab. 1 and 2 summarize PSNR, SSIM, parameter count, and FLOPs for SlowFast-SCI1 (slow-learning backbone: DAUHST) and SlowFast-SCI2 (slow-learning: DPU) against SOTA methods. SlowFast-SCI2 achieves 41.16dB/0.939 on Harvard and 39.79dB/0.963 on ICVL, outperforming all competitors. SlowFast-SCI1 yields 35.81dB/0.867 and 37.68dB/0.950, improving its backbone by 9.71%/7.30% on Harvard and 15.12%/3.26% on ICVL. Both variants use <25% of the parameters and <33% of the FLOPs of their backbones. Compared to SAH-SCI, SlowFast-SCI2 delivers 16.24%/8.06% and 6.79%/1.16% gains on Harvard and ICVL, respectively, with lower resource costs; SlowFast-SCI1 shows similarly strong improvements.
Table 2: Comparisons (PSNR (upper entry ineach cell), SSIM (lower entry in each cell), Params, and FLOPs) between SlowFast-SCI and SOTA methods on 10 simulation scenes of ICVL dataset.
@@ 4.2 Quantitative Results
MST mst CVPR 2022 1.92 25.60
27.07
0.877
31.89
0.909
31.80
0.883
30.43
0.900
35.79
0.931
32.22
0.893
24.05
0.829
27.00
0.842
34.93
0.907
26.30
0.815
30.15
0.879
CST cst ECCV 2022 2.86 36.77
26.61
0.871
31.42
0.904
31.97
0.885
29.36
0.881
36.14
0.935
32.00
0.884
23.26
0.792
26.35
0.833
35.24
0.910
26.72
0.846
29.91
0.874
BIRNAT bisci TPMAI 2023 4.19 1973.66
34.78
0.934
38.41
0.955
36.93
0.951
38.52
0.957
42.17
0.970
37.40
0.945
30.35
0.926
32.59
0.906
37.73
0.932
34.87
0.938
36.37
0.941
RDLUF Dong_2023_CVPR_rdluf CVPR 2023 1.81 115.34
@@
Figure 5: Simulated HSI reconstruction comparisons (middle) on ICVL. The left shows the scene measurement and the right shows the spectral curves corresponding to the selected region. Please zoom in for better view.
@@ 4.3 Qualitative Results
Simulation Data Results. Figs. 4 and 5 depict the simulation HSI reconstruction comparisons between our method and other SOTA methods on Harvard and ICVL datasets, respectively. It can be seen that the other models produce poor reconstruction results with a lot of noise and artifacts, since they lack fast learning on similar data. The results of our SlowFast-SCI in different bands show that the proposed dual speed framework can correct the problem of weak generalization ability of SCI reconstruction due to insufficient training data. This is because our method enhances the generalization of the model to extract more spectral information through self-supervised fast learning and can adapt the model to align it with target scenes at inference. Besides, the spectral density curves show that the reconstruction results of SlowFast-SCI are more similar and correlated with the GT, which proves our method is capable of improving the spectral-dimension consistency of the model.
Real-Scene HSI Reconstruction. We evaluate SlowFast-SCI on real-world HSI using CAVE and KAIST choi2017highkaist , pre-training a two-stage DPU backbone with real CASSI masks under the protocols of Zhang et al. Zhang_2024_CVPR_dpu , Meng et al. meng2020endtsa_cassi , and Cai et al. dauhst . To simulate realistic noise, we inject 11-bit shot noise during both slow- and fast-learning. At test time, raw measurements serve as self-supervised labels for adaptation. As shown in Fig. 6, SlowFast-SCI recovers finer details and suppresses noise more effectively than SOTAs, demonstrating strong real-world potential.
We also provide additional visual comparisons of SlowFast-SCI against SOTAs on scenes from the Harvard, ICVL datasets and real dataset—see the Supplementary Material for details.
Figure 6: Real-scene HSI reconstruction results. Shown are reconstructions of two real-world scenes across 2 selected spectral channels (out of 28), using various state-of-the-art methods. Compared to others, SlowFast-SCI restores finer details and better suppresses noise, highlighting its effectiveness in real-data scenarios.
4.4 Ablation Studies
Fast Unfolding-Stage Configuration Analysis. We assess the effect of different unfolding stages on performance using the Harvard dataset (Tab. 3), where FAST-TTAM denotes the unfolding layer is attached by the adaptation module. Using the 9-stage DPU as a baseline (case a), results from cases (a, c, f) show that both 1-stage and 2-stage variants preserve the key characteristics of the original model. However, cases (d, f) and (e, g) show that the 1-stage variant learns less effectively than the 2-stage network. Thus, despite its faster inference, the 1-stage model is discarded in favor of the better-performing 2-stage variant that enables a balance between reconstruction quality and inference speed.
@@ 4.4 Ablation Studies
Case Stages FAST-TTAM TTA PSNR SSIM Params (M) FLOPs (G)
@@
(c) 1 $\times$ $\times$ 35.67 0.876 0.32 5.39
(d) 1 $\times$ 38.39 0.912 0.35 7.63
(e) 1 38.39 0.903 0.35 7.63
(f) 2 $\times$ $\times$ 35.75 0.876 0.64 10.86
(g) 2 $\times$ 39.97 0.932 0.70 15.17
(h) 2 41.16 0.939 0.70 15.17

### v2

@@ 4.2 Fast Learning
Our slow-learning phase (Sec. 4.1) endows the deep unfolding network with robust physics-informed priors and learnable parameters (i.e., the denoising network). However, when encountering a new spectral domain, naively fine-tuning all unfolding blocks requires nearly the same amount of data and computation as the original training process. Therefore, we aim to develop a more efficient strategy to adapt the model learned during the slow-learning stage.
Eqs. (4) and (5) formally describe the model obtained from the slow-learning phase, where Eq. (4) provides a closed-form solution that is independent of the data distribution. Accordingly, the component of the model that is highly sensitive to the data distribution resides in Eq. (5)—more precisely, in $R(\mathbf{r})$, which is modeled by a learnable denoising network during the slow-learning phase.
By roughly analyzing the statistical roperties as a form of Bayesian inference of this denoising network, we design the Wiener-Filter-Inspired Adapter Modules, which enable the model to adapt efficiently to test samples whose distributions are similar to that of the training data. However, as illustrated in Figs. 4 and 5, the distributions of the two test datasets differ substantially from that of the training dataset. To bridge this distribution gap and further enhance the model’s adaptability, we propose Self-Adaptation-Driven Fast-Learning Unfolding.
4.2.1 Wiener-Filter-Inspired Adapter Modules
How to interpret the statistical properties of general CNNs within the framework of Bayesian inference remains an open question. For clarity of discussion, we denote the optimal denoising filter—i.e., the Wiener filter—for each spectral channel $\lambda_{i}$ corresponding to a training dataset $\mathrm{\mathbf{x}}$ and a specific test sample $\tilde{\mathrm{\mathbf{x}}}$ as follows, respectively:
$\widehat{\mathcal{H}}_{w,\lambda_{i}}[k_{1},k_{2}]=\frac{\mathbb{E}_{\mathbf{x}}\{|\widehat{\mathbf{x}}_{\lambda_{i}}[k_{1},k_{2}]|^{2}\}}{\mathbb{E}_{\mathbf{x}}\{|\widehat{\mathbf{x}}_{\lambda_{i}}[k_{1},k_{2}]|^{2}\}+\mathbb{E}_{\mathbf{\epsilon}}\{|\widehat{\mathbf{\epsilon}}_{\lambda_{i}}[k_{1},k_{2}]|^{2}\}},$ (7)
@@ 4.2.1 Wiener-Filter-Inspired Adapter Modules
Here, $k_{1},k_{2}]$ denotes the spatial frequency components obtained via the discrete Fourier transform (DFT). For simplicity, inter-channel spectral correlations are neglected in this formulation. Detailed derivations of Eqs. (7) and (8) are provided in the supplementary material.
Figure 6: Illustration of four different self-supervised losses used in the fast learning stage. $\Phi$, $Y$, and $\mathcal{F}_{\theta}$ denote the mask, measurement, and model, respectively. $TV(\cdot)$ represents the total variation loss, and RSF stands for random sign-flipping sampling. $T_{g}$ denotes a set of geometric transformations (e.g., shifts, rotations, etc.). The total variation loss $\mathscr{L}_{tv}$ is distribution-independent, serving as a regularization term to preserve image smoothness, whereas $\mathscr{L}_{m}$, $\mathscr{L}_{ei}$, and $\mathscr{L}_{iu}$ are distribution-sensitive.
@@
Table 2: Comparisons (PSNR (upper entry ineach cell), SSIM (lower entry in each cell), Params, and FLOPs) between SlowFast-SCI (SlowFast-SCI1 and SlowFast-SCI2 represent the slow-learning backbone we used were DAUHST and DPU, respectively.) and SOTA methods on 10 simulation scenes of Harvard dataset. Please see the supplementary materials for the complete comparison table.
@@
RDLUF [12] CVPR 2023 1.81 115.34
@@
Figure 7: Simulated HSI reconstruction comparisons (middle) on Harvard (Scene 4) and ICVL (Scene 10). The left shows the scene measurement and the right shows the spectral curves corresponding to the selected region. Please zoom in for better view.
Table 3: Comparisons (PSNR (upper entry ineach cell), SSIM (lower entry in each cell), Params, and FLOPs) between SlowFast-SCI and SOTA methods on 10 simulation scenes of ICVL dataset. Please see the supplementary materials for the complete comparison table.
@@
RDLUF [12] CVPR 2023 1.81 115.34
@@
Figure 8: Real-scene HSI reconstruction results. Shown are reconstructions of two real-world scenes across selected spectral channels (out of 28), using various state-of-the-art methods. Compared to others, SlowFast-SCI restores finer details and better suppresses noise, highlighting its effectiveness in real-data scenarios.
4.2.2 Self-Adaptation-Driven Fast-Learning Unfolding
Adapter-Enhanced Unfolding Network Architecture. Motivated by the analysis in Sec. 4.2.1, we propose an adapter-enhanced unfolding network (AdaptUN). AdaptUN is built upon our previously introduced FastUN (see Sec. 4.1) by appending a lightweight module—namely, the fast test-time adaptation module (FAST-TTAM)—after each unfolding stage. FAST-TTAM jointly leverages spatial and spectral cues to facilitate efficient test-time adaptation. During the adaptation phase, all parameters of the original unfolding backbone are frozen, and only the parameters within FAST-TTAM are learnable. The FAST-TTAM is a lightweight operator with one update step, one parameter $\mu$ estimation network, and two convolution blocks, the spatial convolution block and the spectral convolution block. The structure of parameter $\mu$ estimation network is shown in the upper right of Fig. 2, which is used to estimate the parameter $\mu$. The spatial convolution block consists of two convolution layers with $3\times 3$ filter and one $\mathrm{ReLU}$ function, while the spectral convolution block only includes one convolution layer with $1\times 1$ filter.
Self-Supervised Fine-turning and Test-Time Adaptation. Consider AdaptUN: $\mathcal{F}_{AdaptUN}(\cdot;\omega^{*},\theta)$:
$\mathcal{F}_{AdaptUN}(\cdot;\omega^{*},\theta):\mathrm{\mathbf{y}},\mathbf{\Phi}\rightarrow{\mathrm{\mathbf{x}}},$ (9)
where $\omega^{*}$ denotes the parameters of FastUN frozen in self-supervised fine-turning and test-time adaptation and $\theta$ denotes the parameters of the adapt layers to be updated. Given the same dataset $\{\mathrm{\mathbf{y}}_{i}\}_{i=1}^{N_{d}}$ as in IGFUD, we update $\theta$ by minimizing the following self-supervised loss:
[…]
$\mathscr{L}_{tta}=\mathscr{L}_{m}+\alpha\mathscr{L}_{ei},$ (11)
where $\alpha$ denote the equilibrium constant. To balance efficiency and feasibility, only the two most critical components of the aforementioned self-supervised losses—the measurement consistency loss and the equivariance imaging loss—are utilized during the test time adaptation stage.
5 Experiments
5.1 Experiment Setup
Datasets. The pre-trained model used in slow learning was trained on the CAVE [26] dataset. We adopted two datasets Harvard [9] and ICVL [1] with randomly clipped patches of size $256\times 256\times 28$ for simulation experiments.
Comparison Methods. We compare our SlowFast-SCI on synthetic data and real scenes with state-of-the-art (SOTA) reconstruction methods including TSA-Net [24], MST[5], CST[4], BIRNAT[7], RDLUF[12], SSR-L [43], SAH-SCI [41] , DAUHST [6], and DPU [42].
Implementation Details. The implementation details of SlowFast-SCI are given in the supplementary materials.
5.2 Quantitative Results
Comparison on Harvard and ICVL.
Tab. 2 and 3 summarize PSNR, SSIM, Params, and FLOPs for SlowFast-SCI1 (slow-learning backbone: DAUHST) and SlowFast-SCI2 (slow-learning backbone: DPU) against SOTA methods. SlowFast-SCI2 achieves 41.16dB/0.939 on Harvard and 39.79dB/0.963 on ICVL, outperforming all competitors. SlowFast-SCI1 yields 35.81dB/0.867 and 37.68dB/0.950, improving its backbone by 9.71%/7.30% on Harvard and 15.12%/3.26% on ICVL. Both variants use <<25% of the Params and <<33% of the FLOPs of their backbones. Compared to SAH-SCI, SlowFast-SCI2 achieves 16.24%/8.06% and 6.79%/1.16% gains on Harvard and ICVL, respectively, with lower resource costs; SlowFast-SCI1 attains similar improvements.
@@ 5.3 Qualitative Results
Simulation Data Results. Fig. 7 epicts the simulation HSI reconstruction comparisons between our method and other SOTA methods on Harvard and ICVL datasets, respectively. It can be seen that the other models produce poor reconstruction results with a lot of noise and artifacts, since they lack fast learning on similar data. The results of our SlowFast-SCI in different bands show that the proposed dual speed framework can correct the problem of weak generalization ability of SCI reconstruction due to insufficient training data. This is because our method enhances the generalization of the model to extract more spectral information through self-supervised fast learning and can adapt the model to align it with target scenes at inference. Besides, the spectral density curves show that the reconstruction results of SlowFast-SCI are more similar and correlated with the GT, which proves our method is capable of improving the spectral-dimension consistency of the model.
Real-Scene HSI Reconstruction. We evaluate SlowFast-SCI on real-world HSI using CAVE and KAIST [10], pre-training a two-stage DPU backbone with real CASSI masks under the protocols of Zhang et al. [42], Meng et al. [24], and Cai et al. [6]. To simulate realistic noise, we inject 11-bit shot noise during both slow- and fast-learning. At test time, raw measurements serve as self-supervised labels for adaptation. As shown in Fig. 8, SlowFast-SCI recovers finer details and suppresses noise more effectively than SOTAs, demonstrating strong real-world potential. We also provide additional visual comparisons of SlowFast-SCI against SOTAs on scenes from the Harvard, ICVL datasets and real dataset—see the Supplementary Material for details.
@@
Case Stages FAST-TTAM TTA PSNR SSIM Params (M) FLOPs (G) Time (s)
@@
(c) 9 33.75 0.825 3.16 67.92 $\sim$45.9
(d) 1 $\times$ $\times$ 35.67 0.876 0.32 5.39
(e) 1 $\times$ 38.39 0.912 0.35 7.63
(f) 1 38.39 0.903 0.35 7.63
(g) 2 $\times$ $\times$ 35.75 0.876 0.64 10.86
(h) 2 $\times$ 39.97 0.932 0.70 15.17
(i) 2 41.16 0.939 0.70 15.17 $\sim$10.9
5.4 Ablation Studies
Fast Unfolding-Stage Configuration Analysis. We assess the effect of different unfolding stages on performance using the Harvard dataset (Tab. 4), where FAST-TTAM denotes the unfolding layer is attached by the adaptation module. Using the 9-stage DPU as a baseline (case a), results from cases (a, d, g) show that both 1-stage and 2-stage variants preserve the key characteristics of the original model. However, cases (e, h) and (f, i) show that the 1-stage variant learns less effectively than the 2-stage network. Thus, despite its faster inference, the 1-stage model is discarded in favor of the better-performing 2-stage variant that enables a balance between reconstruction quality and inference speed. In particular, case (c) can be regarded as the implementation of AdaptNet [46] under the SCI configuration. Notably, AdaptNet fails to adapt to OOD data and suffers from slow inference. In contrast, case (i) demonstrates that our proposed method not only achieves effective test-time alignment with OOD test samples but also significantly accelerates inference.
Fast Learning Ablation. We analyze the performance of FAST-TTAM and TTA in the fast learning. Tab. 4 (d, e) and (g, h) show that the proposed FAST-TTAM can effectively transfer the fast unfolding variant to the target domain through self-supervised fine-turning. Besides, case (h, i) in Tab. 4 shows self-supervised test-time adaptation can further align the fast unfolding variant with target scenes.